6. An individual is willing to forgo increasing amounts of good Y for each additional unit of good X. Which of the following utility functions is consistent with these preferences? a. U(X, Y) =X+y? b. U(X, Y) (X1/2Y1/2)2 c. U(X,Y) Y2 x2
Answer #1

6. An individual is willing to forgo increasing amounts of good Y for each additional units of good X. In such situation

a.  U(X,Y) = X + Y2 is the utility functions which is consistent with these preferences.

Explanation

U(X,Y) = X + Y2 is the only utility function in which marginal utility of good X is increases. where as in the utility function b,c and d, marginal utility of good X is decreases. Therefore, only   a.  U(X,Y) = X + Y2 is the utility functions which is consistent with individual who is willing to forgo increasing amounts of good Y for each additional units of good X.